### v1

# Simulated Interactive Debugging

Abstract.
Debugging software, i.e., the localization of faults and their repair, is a main activity in software engineering. Therefore, effective and efficient debugging is one of the core skills a software engineer must develop. However, the teaching of debugging techniques is usually very limited or only taught in indirect ways, e.g., during software projects. As a result, most Computer Science (CS) students learn debugging only in an ad-hoc and unstructured way. In this work, we present our approach called Simulated Interactive Debugging that interactively guides students along the debugging process. The guidance aims to empower the students to repair their solutions and have a proper ”learning” experience. We envision that such guided debugging techniques can be integrated into programming courses early in the CS education curriculum. To perform an initial evaluation, we developed a prototypical implementation using traditional fault localization techniques and large language models. Students can use features like the automated setting of breakpoints or an interactive chatbot. We designed and executed a controlled experiment that included this IDE-integrated tooling with eight undergraduate CS students. Based on the responses, we conclude that the participants liked the systematic guidance by the assisted debugger. In particular, they rated the automated setting of breakpoints as the most effective, followed by the interactive debugging and chatting, and the explanations for how breakpoints were set. In our future work, we will improve our concept and implementation, add new features, and perform more intensive user studies.
Intelligent Tutoring, Debugging, Software Engineering, Education
1. Introduction
Despite the paradigm shift towards Artificial Intelligence (AI)-assisted software development, particularly with the emergence of Large Language Models (LLMs), there remains a high demand for qualified software engineers capable of developing reliable, high-quality software. With more auto-generated code, it will become even more critical that junior software developers can effectively debug software and solve bugs, even in code they have not written themselves. Radermacher et al. (Radermacher et al., 2014) explored knowledge deficiencies of graduate students from an industry perspective. One of the most frequently mentioned issues with software tools has been version control systems and debuggers. While we have seen the integration of version control systems like Git in the Computer Science (CS) and Software Engineering (SE) curriculum, e.g., with the usage of GitHub Classroom, we have not seen much innovation regarding debugging education. Michaeli and Romeike (Michaeli and Romeike, 2019) noted that there are only few studies investigating the ”explicit teaching of debugging” and students are often left alone to learn debugging ”the hard way”. Our experience confirms this observation: most students still learn (interactive) debugging of code in an ad-hoc and unstructured way by trial and error. Using ”print/log” statements is often the only extent of students’ experience in debugging. They do not receive proper guidance in efficient program comprehension and how to debug programs effectively. Current CS/SE research efforts focus on general programming education, which helps to provide scalable alternatives to mentor and guide a rising number of students. Automated Program Repair (APR) techniques (Gulwani et al., 2018; Hu et al., 2019; Zhang et al., 2024) can help to produce patches for incorrect submissions, which then can be used as a basis for personalized feedback, or even automated grading (Fan et al., 2023). In fact, we already have seen the application of such techniques in the CS/SE curriculum focusing on repairing students’ solutions and providing feedback (Kazemitabaar et al., 2024b; Fan et al., 2024). However, the existing related work does not address actual debugging.
Our research objective is to understand students’ needs in learning to debug and provide automated means to guide them effectively along the debugging process.
As part of this objective, we strive to seek answers to questions like ”How can we guide the debugging process using artifacts generated from APR/SE techniques?” and ”How can LLMs be integrated into the debugging process for the benefit of the student?”
As a first step to achieve our research objective, we propose the concept of Simulated Interactive Debugging. The essence of this concept is to create a simulation of an interactive debugging experience; ”simulated” because the solution is known in the education context due to given reference implementations and artifacts like fault locations and patches that can be auto-generated with SE and APR techniques. Our goal is to provide a controlled but supportive environment, in which the students debug the problem on their own. Similar to a human tutor who would not directly reveal the solution and instead help the student to understand and fix the issue in their code. Our vision of such an assisted debugger requires integration into coding tools like Integrated Development Environments (IDEs) to guide students actively during the debugging process. Based on APR, fault localization results, and the usage of LLMs, we can show potential program problematic locations, explain the problem at these locations, set breakpoints automatically, identify relevant variable values, highlight them, and provide partial or complete hints on how to repair the student’s program.
As part of a design science methodology (Hevner et al., 2004; Wieringa, 2014), we developed a prototype of the proposed concept as a Visual Studio Code IDE extension. We conducted a pilot user study with eight undergraduate CS students. The results show that students have little experience with (interactive) debugging but enjoy the provided guidance, notably the automatic breakpoint setting and the interaction with the AI-based chatbot. In summary, our core contributions are:
  • the concept of simulated interactive debugging to guide CS students along the interactive debugging process
  • the implementation of this concept as an IDE plugin
  • a pilot user study to show the practicality and usability of our concept and its implementation
2. Related Work
Intelligent Tutoring Systems
Due to the rise in the number of CS students (Masters, 2011), educators need help to cope with the number of assignment submissions and the needed feedback. APR techniques are currently being explored to supplement the missing human tutor, e.g., by realizing so-called Intelligent Tutoring Systems (ITS). Early works in this domain attempt to repair student submissions to generate feedback (Gulwani et al., 2018; Hu et al., 2019) and automatically grade programming assignments (Fan et al., 2023; Birillo et al., 2022). More recently, LLM-based approaches have been explored intensively to generate personalized feedback and hints for programming assignments (Zhang et al., 2024; Liffiton et al., 2024; Hou et al., 2024; Zhao et al., 2024; S Kumar et al., 2024). Several experiences have been reported where LLM-based programming assistants have been deployed in the classroom context (Abolnejadian et al., 2024; Kazemitabaar et al., 2024b; Fan et al., 2024; Kurniawan et al., 2023). LLMs also can be combined with traditional techniques like static analysis to provide incremental hints for programming tasks (Birillo et al., 2024a). Moreover, existing works (Koutcheme et al., 2024; Phung et al., 2023) benchmarked various models regarding their capabilities to support programming education. Overall, the existing works focus on repairing programming assignments and generating student feedback; they do not address the teaching of interactive debugging.
Debugging Education
To get an overview of the teaching of debugging, we refer to McCauley et al. (McCauley et al., 2008), who conducted a systematic literature review. Michaeli and Romeike (Michaeli and Romeike, 2019) recently explored the influence of teaching systematic debugging concepts with an intervention study concluding that explicitly teaching debugging skills positively affects debugging self-efficacy. Recent LLM-inspired, debugging-related research focuses on Socratic questioning (Al-Hossami et al., 2024) and generating so-called debugging quizzes (Phung et al., 2023; Padurean et al., 2024) that ask students to reason about a given buggy program and design bug-revealing test cases. While the reported experiences and the existing approaches and tools can help students build general debugging skills, to the best of our knowledge, no approach attempts to guide students through the interactive debugging process.
AI in CS Education
At the latest, with the emergence of LLMs, AI also arrived in the domain of CS education. Shein (Shein, 2024) argues that while we still need to teach students the fundamentals of programming, such teaching will become more ”alive” through AI, e.g., via chatbots. Denny et al. (Denny et al., 2024) investigated desirable characteristics for AI teaching assistants and concluded that students enjoy AI-based learning support and actually preferred tailored scaffolding instead of mere responses and the revealing of solutions. Kazemitabaar et al. (Kazemitabaar et al., 2024a) explored the design space for such AI tools and highlighted the importance of step-by-step guidance and an interactive dialog with the AI. Following the insights of these previous works, our proposed concept for simulated interactive debugging guides the student in understanding the problem and identifying the programming faults instead of just providing feedback on how to fix it. Further, we also offer an AI-based chatbot with which the students can engage in a controlled environment.
3. Concept: Simulated Interactive Debugging
To tackle the shortcomings of CS education in program debugging, we propose the following concept. Using technology from automated software engineering (SE), in particular from automated program repair (APR), we can identify the issue in a student’s buggy program and can generate a corresponding solution. Note that such a solution can structurally and syntactically differ from the reference implementation provided by a lecturer. Now, knowing the fault and where to fix it, we can use this knowledge to guide the student not only in fixing the issue, but step-by-step in understanding the program’s faults, identifying the potential fix locations, and eventually fixing the problem—similar to how a human tutor would guide the student along the debugging process. Therefore, we designed and developed an AI-assisted debugger that offers students a simulated interactive debugging experience. Students receive help in running tests, using the interactive debugger in identifying the underlying fault in the problem, and in fixing the actual problem.
The Role of LLMs
This concept fits into a larger research endeavor of ours, where we aim to develop an active learning environment integrating APR through LLMs and other SE tools with interactive debugging. This concept is designed to provide interactive guidance to students, thereby enhancing their learning experience in programming. Furthermore, our concept exposes students to AI tools, in particular LLMs, in a guided and controlled way. LLMs can be deployed in various roles in this concept, e.g., to generate hints and explanations in natural language, as well as core technology for an interactive chatbot. Based on the generated artifacts from techniques like spectrum-based fault localization (SBFL) (Wong et al., 2016), static code analysis, symbolic execution, code synthesis and LLM prompting, we can set breakpoints automatically, identify relevant variable values, and highlight them for the students. Such a process can also be supported by Socratic questioning (Al-Hossami et al., 2024).
Workflow
The concept has two types of users: the students and the lecturers, although the main focus is on the students. Both user groups provide inputs and receive feedback. The lecturer provides the programming assignment, the reference implementation, and a test suite consisting of private and public test cases. The student provides a buggy solution and (optionally) a set of custom test cases. There are three different use cases: (1) the primary use case is the assisted debugging (discussed in Section 3.1), (2) the test assessment for lecturers, and (3) the test assessment for custom tests from the student. The test assessment for lecturers checks whether the given test suite is strong enough to catch errors and perform effective fault localization. This can be realized via mutation testing, and additional tests could be generated, e.g., via evolutionary test generation (Fraser and Arcuri, 2011). By offering additional features for the lecturer, our goal is also to increase confidence in the overall system, which is crucial for its successful deployment in programming courses. The test assessment for customs tests from the student works with the idea that when the student submits custom tests, these can be executed against the reference implementation from the lecturer. If they fail on the reference implementation, they are flagged and can also be corrected (because the correct output for the inputs of such test cases could be extracted via the reference implementation). To help students, we also can propose new test cases to them, e.g., to explore certain boundary cases or to increase overall code coverage.
Deployment
Birillo et al. (Birillo et al., 2024b) already argued that combining programming education with the usage of IDEs helps to familiarize students with industrial technologies. Building on top of this idea, our concept is envisioned to be developed as an IDE plugin.
3.1. Assisted Debugging
Figure 1. Envisioned dialogue between student and AI bot.
Figure 1 showcases the envisioned interaction between the student and the AI assistant. First, the AI assistant can (1) clarify/explain failing test cases and the observed failures. Next, it can (2) automatically set breakpoints at locations of interest. These locations can be identified, e.g., by using SBFL. Setting such breakpoints enables the unexperienced student to use the interactive debugger to step through the test execution. Based on LLMs, we also can (3) generate hints in natural language explaining why these breakpoints are of interest. In cases where a student’s understanding of the problem is far from what is expected, the AI assistant can (4) deliver incremental guidance, e.g., via progressive hints, to ensure that the student understands the problem correctly. If the student is still stuck, the AI assistant can (5) partially reveal why a certain test case execution fails. Furthermore, we can (6) help the student in the interactive debugging by identifying and highlighting interesting program variables that the student can observe during the debugging session. This interaction aims to guide the student to repair their own solution and to have a proper ”learning” experience.
Figure 2. Screenshot from our prototypical implementation. The left part shows the code and the automatically set breakpoints. The right part shows the feature selection and the dialogue with the AI-based chatbot.
Figure 3. Participants’ programming/debugging experience.
Figure 4. Participants’ assessment of the provided features.
4. Pilot User Study
To provide first insights towards our research goals, we executed a pilot user study. We investigate the practicality of our assisted debugging concept, as well as the usability of our implementation.
4.1. Implementation
As a study prototype, we realized the Simulated Interactive Debugging concept as a Visual Studio Code IDE extension. Our concept is language-agnostic; however, due to the programming courses at our institution, we decided to first focus on supporting Python. It provides two core features for the assisted debugger: the automatic setting of breakpoints (based on SBFL with FauxPy (Rezaalipour and Furia, 2024)) and a chat interface providing tailored debugging hints. Each feature can be triggered via a VSCode command. The chat interface connects the participant with an LLM-powered chatbot. The user can choose from two modes: the Generate Hints mode, which only allows the student to get step-by-step hints, and the Interactive Debugging Guidance mode, which allows the student to engage in a dialogue with the LLM in a controlled way. The student can discuss code issues and query about how to proceed with debugging or improving the code. Figure 2 shows an annotated screenshot of our implementation. As underlying LLM, we used OpenAI gpt-3.5-turbo. Note that in the implementation of our current prototype, we omitted the support for the lecturer’s and the student’s test assessment because our primary interest was in the feedback for the actual debugging.
4.2. Study Setup
Experiment Structure
The study started with a short briefing, the IRB signing, and the completion of a profiling survey. Then, we spent 15 minutes for providing an introduction to the topic of debugging and for demonstrating the features of our tool. Afterwards, the participants had 40 minutes to solve two coding tasks. The tasks originate from LeetCode and represent programming tasks of medium difficulty that should not overwhelm or underwhelm the students. Both were provided with pre-existing buggy solutions, along with corresponding passing and failing test cases. In the first task, named maxOperations (lee, 2025b), the participants have to fix an algorithm that is supposed to calculate the maximum number of times, to pick two numbers from a given array whose sum equals the parameter $k$. In the second task, named longestOnes (lee, 2025a), the participants have to fix the algorithm that is supposed to compute the maximum number of consecutive 1’s in a binary array if one can flip at most $k$ 0’s. The participants had first to understand the task’s description, then find the errors in the provided solution, and finally provide a modified implementation that passes all test cases. To remember the available assisted debugging features, we also provided each participant with a handout describing a possible workflow to solve the tasks (see replication artifact [^1] Please contact the corresponding author to request the replication artifact.). We informed the participants that the chatbot’s hints might not be entirely accurate. In the end, the participants were asked to complete a usability questionnaire (Brooke, 1996) and a post-task survey. The study was timed for 70 minutes, and the participants were compensated with 30 SGD.
Participants
For our pilot study, we recruited eight first-year undergraduate CS students. Based on our profiling survey results, all of them are beginners in the field of programming, and hence, are in our target group. Figure 3 illustrates the self-classified experience levels of our participants: Most of them have programming experience of less than two years and can only write basic Python programs. 6/8 participants mentioned that they can debug basic errors but struggle with complex ones. At the same time, 6/8 have never heard of or never used an interactive debugger. Those who understand the concept of an interactive debugger still prefer using print/log statements to debug their code.
4.3. Results & Discussion
In the following, we discuss the findings from our user study regarding the feedback on the existing features of the assisted debugger, the overall usability of our prototype, the participant’s confidence of manually setting breakpoints, and additionally desired features.
Existing Features
In total, all eight participants were able to solve the two programming tasks in the provided time. As shown in Figure 4, the most effective feature is the automatic breakpoint setting (5/8 agreement), followed by interactive debugging and chatting (4/8) and the explanation of automatic breakpoints (4/8). The features for progressive hints and test error explanation are only positively evaluated by 2/8 students. The progressive hints helped the students to identify where their code has a fault but did not reveal the solution. Instead, it attempted to guide the students to find the solution on their own to foster a learning effect. However, the participants perceived this feature as often too vague or too general. For most participants, the test error explanation helped to analyze the coding issue, while the feature itself could be made more interactively, e.g., allowing for different/alternative explanations. The explanation of the automatically set breakpoints helped most students to understand the wrong parts. Therefore, we conclude that it is a good addition to the automatic setting of breakpoints. The interactive debugging and chatting was able to add more detailed information to the debugging efforts. One participant mentioned that this even accelerated the overall debugging process. Others still found that the responses were vague and that more prompts could improve the feature. The automatic setting breakpoints feature was considered very effective as it set the breakpoints at the right places. One participant summarized its effectiveness as follows:
”It did get the breakpoint in the right place, which was a good start in debugging.”
One of the first hurdles of conventional interactive debugging is to set breakpoints. Our concept of simulated interactive debugging helps them to get started, and therefore, guides them along a deeper understanding of their coding errors and strives to develop debugging skills. The following quote from one of our participants about the interactive debugging and chatting highlights another aspect:
”It helped me brainstorm while guiding me and
not letting me fully rely on AI”
Our goal is to guide students to learn debugging but also the usage of AI-based tools like ChatGPT. By constructing a chatbot that uses LLMs, we can control the prompts and the contents of responses, and hence, still foster a learning experience for the students, e.g., by filtering direct solutions.
Usability
Using the standardized system usability survey by Brooke (Brooke, 1996), we can conclude that overall, the participants are satisfied with the prototype’s usability. The average System Usability Scale (SUS) is around 65 (out of 100), providing a good basis for our future work. For example, 7/8 participants would like to use the system frequently, and 5/8 agreed that the system is easy to use. These findings confirm that our efforts are heading in the right direction but also leave space for further improvement.
Debugging Process
After completing the tasks, we also asked the participants how confident they are about setting breakpoints on their own. The responses indicated mixed confidence with a tendency to lower confidence: 1/8 participants strongly agreed to be confident, 2/8 agreed, 2/8 noted a neutral opinion, and 3/8 disagreed. Afterwards, we asked them to set breakpoints for a given code snippet (see Listing 1). The interesting lines are 6 to 7 because the handling of None values is missing.
Listing 1: Code snippet for setting a manual breakpoints.
1# Calculate the average of a list of
2# grades, if None, you should continue
3# to the next grade
4def calculate_average (grades):
5 total = 0
6 for grade in grades:
7 total += grade
8 average = total / len (grades)
9 return average
10
11# Test the function
12grades = [85, 90, 78, None, 92, None]
13average_grade = calculate_average(grades)
14print(f”Theaverageis:{average_grade}”)
6/8 participants would set breakpoints either at lines 6 or 7, or both, which shows that actually most of the participants have a correct intuition about where to set the breakpoint for further inspection. While we see good results for this relatively simple example, we conclude that due to overall low confidence, the automated setting of breakpoints would still be an interesting feature. However, such a feature could be applied only as hints or as confirmation for the students’ manually set breakpoints. Additionally, we asked the participants about their perceived key difference to their usual debugging process. The majority (6/8) mentioned that our simulated interactive debugging tool helped them to debug the coding issue and guided them systematically. However, one participant mentioned that interactive debugging could be more time-consuming. The following two quotes are representative responses:
”Normally I would just trial and error until I got it right, but this tool allowed me to systematically identify the problems in the code step by step.” ”My usual debugging process includes using print statements and doing everything myself but sometimes I just get tired and ask ChatGPT. Simulated Interactive Debugging helped me think while guiding me through the right process.”
Additional Features
At the end of the post-task survey, we asked the participants about any additional features they would like to see in the Simulated Interactive Debugger tool that have not been implemented yet. 6/8 participants indicated that further hints about what and how to change would be helpful. For example, when the students really struggle, the chatbot could provide more detailed explanations and propose changes that the student could select. Additionally, one participant mentioned that the user interface could be improved to enable easier switching between the debugging information from the assisted debugger and the actual interactive debugging tool in the IDE. Lastly, one participant highlighted that integrating some automated test case generation would be helpful. For example, such test cases could be provided if the students cannot find an appropriate input for debugging.
4.4. Threats to Validity
Since we only performed a first pilot study with eight CS students, our results may not generalize. We focused our user study on a small number of students to receive focused feedback for the next development phase. Nevertheless, the results of our profiling survey showed that the participants are in our target, i.e., beginners in programming with no or only very limited experience with (interactive) debugging. For the next cycle of user studies, we will attempt to reach a larger set of participants. In order to mitigate the threat of manual errors in our analysis, we ensured that all results were analyzed and agreed upon by two authors independently.
5. Conclusion and Future Work
We proposed the concept of simulated interactive debugging as a first step to automate the teaching of debugging skills. Incorporating such techniques in the CS curriculum will be essential to teaching debugging beyond ad-hoc approaches and guiding students in using AI-based programming tools. In future, we plan to improve our interactive chatbot with more guidance, e.g., by using a state machine-driven approach, similar to Bouzenia et al. (Bouzenia et al., 2024). We also will explore the integration of several state-of-the-art APR techniques, implement the test assessment features for lecturers and students, and investigate how we can automatically help students to generate and improve their test suites.

### v2

@@
 Simulated Interactive Debugging
-Abstract.
-Debugging software, i.e., the localization of faults and their repair, is a main activity in software engineering.
+Abstract
+Debugging software, i.e., the localization of faults and their repair, is a key activity in software engineering.
 Therefore, effective and efficient debugging is one of the core skills a software engineer must develop.
@@
 In this work, we present our approach called Simulated Interactive Debugging that interactively guides students along the debugging process.
-The guidance aims to empower the students to repair their solutions and have a proper ”learning” experience.
+The guidance aims to empower the students to repair their solutions and have a proper learning experience.
 We envision that such guided debugging techniques can be integrated into programming courses early in the CS education curriculum.
-To perform an initial evaluation, we developed a prototypical implementation using traditional fault localization techniques and large language models.
+We developed a prototypical implementation using traditional fault localization techniques and large language models.
 Students can use features like the automated setting of breakpoints or an interactive chatbot.
-We designed and executed a controlled experiment that included this IDE-integrated tooling with eight undergraduate CS students.
-Based on the responses, we conclude that the participants liked the systematic guidance by the assisted debugger.
-In particular, they rated the automated setting of breakpoints as the most effective, followed by the interactive debugging and chatting, and the explanations for how breakpoints were set.
-In our future work, we will improve our concept and implementation, add new features, and perform more intensive user studies.
+We designed and executed a small-scale, controlled experiment with eight undergraduate CS students.
+Based on the responses, we conclude that the participants liked the systematic guidance.
+They rated the automated setting of breakpoints as most effective, followed by the interactive debugging and chatting, and the breakpoint explanations.
+In future, we will extend our concept and implementation, and perform more intensive user studies.
+Index Terms:
 Intelligent Tutoring, Debugging, Software Engineering, Education
-1. Introduction
+I Introduction
 Despite the paradigm shift towards Artificial Intelligence (AI)-assisted software development, particularly with the emergence of Large Language Models (LLMs), there remains a high demand for qualified software engineers capable of developing reliable, high-quality software.
@@
 Radermacher et al.
-(Radermacher et al., 2014) explored knowledge deficiencies of graduate students from an industry perspective.
+[1] explored knowledge deficiencies of graduate students from an industry perspective.
 One of the most frequently mentioned issues with software tools has been version control systems and debuggers.
 While we have seen the integration of version control systems like Git in the Computer Science (CS) and Software Engineering (SE) curriculum, e.g., with the usage of GitHub Classroom, we have not seen much innovation regarding debugging education.
-Michaeli and Romeike (Michaeli and Romeike, 2019) noted that there are only few studies investigating the ”explicit teaching of debugging” and students are often left alone to learn debugging ”the hard way”.
+Michaeli and Romeike [2] noted that there are only a few studies investigating the “explicit teaching of debugging” and students are often left alone to learn debugging “the hard way”.
 Our experience confirms this observation: most students still learn (interactive) debugging of code in an ad-hoc and unstructured way by trial and error.
-Using ”print/log” statements is often the only extent of students’ experience in debugging.
+Using “print/log” statements is often the only extent of students’ experience in debugging.
 They do not receive proper guidance in efficient program comprehension and how to debug programs effectively.
 Current CS/SE research efforts focus on general programming education, which helps to provide scalable alternatives to mentor and guide a rising number of students.
-Automated Program Repair (APR) techniques (Gulwani et al., 2018; Hu et al., 2019; Zhang et al., 2024) can help to produce patches for incorrect submissions, which then can be used as a basis for personalized feedback, or even automated grading (Fan et al., 2023).
-In fact, we already have seen the application of such techniques in the CS/SE curriculum focusing on repairing students’ solutions and providing feedback (Kazemitabaar et al., 2024b; Fan et al., 2024).
-However, the existing related work does not address actual debugging.
-Our research objective is to understand students’ needs in learning to debug and provide automated means to guide them effectively along the debugging process.
-As part of this objective, we strive to seek answers to questions like ”How can we guide the debugging process using artifacts generated from APR/SE techniques?” and ”How can LLMs be integrated into the debugging process for the benefit of the student?”
-As a first step to achieve our research objective, we propose the concept of Simulated Interactive Debugging.
-The essence of this concept is to create a simulation of an interactive debugging experience; ”simulated” because the solution is known in the education context due to given reference implementations and artifacts like fault locations and patches that can be auto-generated with SE and APR techniques.
+Automated Program Repair (APR) techniques [3, 4, 5] can help to produce patches for incorrect submissions, which then can be used as a basis for personalized feedback, or even automated grading [6].
+In fact, we already have seen the application of such techniques in the CS/SE curriculum focusing on repairing students’ solutions and providing feedback [7, 8].
+However, the existing related work in CS education does not address actual debugging.
+Therefore, our research objective is to understand students’ needs in learning to debug and provide automated means to guide them effectively along the debugging process.
+As a first step to achieve our research objective, we propose Simulated Interactive Debugging.
+The essence of this concept is to create a simulation of an interactive debugging experience; “simulated” because the solution is known in the education context due to given reference implementations and auto-generated artifacts like fault locations and patches.
 Our goal is to provide a controlled but supportive environment, in which the students debug the problem on their own.
-Similar to a human tutor who would not directly reveal the solution and instead help the student to understand and fix the issue in their code.
-Our vision of such an assisted debugger requires integration into coding tools like Integrated Development Environments (IDEs) to guide students actively during the debugging process.
-Based on APR, fault localization results, and the usage of LLMs, we can show potential program problematic locations, explain the problem at these locations, set breakpoints automatically, identify relevant variable values, highlight them, and provide partial or complete hints on how to repair the student’s program.
-As part of a design science methodology (Hevner et al., 2004; Wieringa, 2014), we developed a prototype of the proposed concept as a Visual Studio Code IDE extension.
-We conducted a pilot user study with eight undergraduate CS students.
-The results show that students have little experience with (interactive) debugging but enjoy the provided guidance, notably the automatic breakpoint setting and the interaction with the AI-based chatbot.
+Similar to a human tutor, we would not directly reveal the solution and instead help the student to understand and fix the issue.
+Our vision of such an assisted debugger requires integration into coding tools like Integrated Development Environments (IDEs) to actively guide students.
+Based on APR, fault localization, and LLMs, we can show potential problematic locations, explain the issues at these locations, set breakpoints, identify relevant variable values, and provide repair hints.
+As part of a design science methodology [9, 10], we developed a prototype as a VS Code IDE extension and conducted a pilot user study with eight undergraduate CS students.
+The results indicate that students have little experience with (interactive) debugging but enjoy the guidance, notably the automatic breakpoint setting and the interaction with the chatbot.
 In summary, our core contributions are:
@@
   • a pilot user study to show the practicality and usability of our concept and its implementation
-2. Related Work
+II Related Work
 Intelligent Tutoring Systems
-Due to the rise in the number of CS students (Masters, 2011), educators need help to cope with the number of assignment submissions and the needed feedback.
+Due to the rise in the number of CS students [11], educators need help to cope with the number of assignment submissions.
 APR techniques are currently being explored to supplement the missing human tutor, e.g., by realizing so-called Intelligent Tutoring Systems (ITS).
-Early works in this domain attempt to repair student submissions to generate feedback (Gulwani et al., 2018; Hu et al., 2019) and automatically grade programming assignments (Fan et al., 2023; Birillo et al., 2022).
-More recently, LLM-based approaches have been explored intensively to generate personalized feedback and hints for programming assignments (Zhang et al., 2024; Liffiton et al., 2024; Hou et al., 2024; Zhao et al., 2024; S Kumar et al., 2024).
-Several experiences have been reported where LLM-based programming assistants have been deployed in the classroom context (Abolnejadian et al., 2024; Kazemitabaar et al., 2024b; Fan et al., 2024; Kurniawan et al., 2023).
-LLMs also can be combined with traditional techniques like static analysis to provide incremental hints for programming tasks (Birillo et al., 2024a).
-Moreover, existing works (Koutcheme et al., 2024; Phung et al., 2023) benchmarked various models regarding their capabilities to support programming education.
-Overall, the existing works focus on repairing programming assignments and generating student feedback; they do not address the teaching of interactive debugging.
+Early works in this domain attempt to repair student submissions to generate feedback [3, 4] and automatically grade programming assignments [6, 12].
+More recently, LLM-based approaches have been explored to generate personalized feedback and hints for programming assignments [5, 13, 14, 15, 16].
+Several experiences have been reported where LLM-based programming assistants have been deployed in the classroom context [7, 8, 17, 18].
+LLMs also can be combined with traditional techniques like static analysis to provide incremental hints for programming tasks [19].
+Moreover, existing works [20, 21] benchmarked various models regarding their capabilities to support programming education.
+Overall, the existing works focus on repairing programming assignments and generating feedback; they do not address the teaching of interactive debugging.
 Debugging Education
 To get an overview of the teaching of debugging, we refer to McCauley et al.
-(McCauley et al., 2008), who conducted a systematic literature review.
-Michaeli and Romeike (Michaeli and Romeike, 2019) recently explored the influence of teaching systematic debugging concepts with an intervention study concluding that explicitly teaching debugging skills positively affects debugging self-efficacy.
-Recent LLM-inspired, debugging-related research focuses on Socratic questioning (Al-Hossami et al., 2024) and generating so-called debugging quizzes (Phung et al., 2023; Padurean et al., 2024) that ask students to reason about a given buggy program and design bug-revealing test cases.
-While the reported experiences and the existing approaches and tools can help students build general debugging skills, to the best of our knowledge, no approach attempts to guide students through the interactive debugging process.
+[22], who conducted a systematic literature review.
+Michaeli and Romeike [2] recently explored the influence of teaching systematic debugging concepts with an intervention study concluding that explicitly teaching debugging skills positively affects debugging self-efficacy.
+Recent LLM-inspired, debugging-related research focuses on Socratic questioning [23] and generating so-called debugging quizzes [21, 24] that ask students to reason about a given buggy program and design bug-revealing test cases.
+While the reported approaches and tools can help students build general debugging skills, to the best of our knowledge, no approach guides through the interactive debugging process.
 AI in CS Education
-At the latest, with the emergence of LLMs, AI also arrived in the domain of CS education.
-Shein (Shein, 2024) argues that while we still need to teach students the fundamentals of programming, such teaching will become more ”alive” through AI, e.g., via chatbots.
+Most recently, with the emergence of LLMs, AI has arrived in the domain of CS education.
+Shein [25] argues that while we still need to teach students the fundamentals of programming, such teaching will become more “alive” through AI, e.g., via chatbots.
 Denny et al.
-(Denny et al., 2024) investigated desirable characteristics for AI teaching assistants and concluded that students enjoy AI-based learning support and actually preferred tailored scaffolding instead of mere responses and the revealing of solutions.
+[26] investigated desirable characteristics for AI teaching assistants and concluded that students enjoy AI-based learning support and actually preferred tailored scaffolding instead of mere responses and the revealing of solutions.
 Kazemitabaar et al.
-(Kazemitabaar et al., 2024a) explored the design space for such AI tools and highlighted the importance of step-by-step guidance and an interactive dialog with the AI.
-Following the insights of these previous works, our proposed concept for simulated interactive debugging guides the student in understanding the problem and identifying the programming faults instead of just providing feedback on how to fix it.
+[27] explored the design space for such AI tools and highlighted the importance of step-by-step guidance and an interactive dialog with the AI.
+Following the insights of these works, our proposed concept for Simulated Interactive Debugging guides the student in understanding the problem and identifying the programming faults instead of just providing feedback on how to fix it.
 Further, we also offer an AI-based chatbot with which the students can engage in a controlled environment.
-3. Concept: Simulated Interactive Debugging
+III Concept: Simulated Interactive Debugging
 To tackle the shortcomings of CS education in program debugging, we propose the following concept.
-Using technology from automated software engineering (SE), in particular from automated program repair (APR), we can identify the issue in a student’s buggy program and can generate a corresponding solution.
+Using technology from automated SE, in particular from APR, we can identify the issue in a student’s buggy program and can generate a corresponding solution.
 Note that such a solution can structurally and syntactically differ from the reference implementation provided by a lecturer.
-Now, knowing the fault and where to fix it, we can use this knowledge to guide the student not only in fixing the issue, but step-by-step in understanding the program’s faults, identifying the potential fix locations, and eventually fixing the problem—similar to how a human tutor would guide the student along the debugging process.
-Therefore, we designed and developed an AI-assisted debugger that offers students a simulated interactive debugging experience.
+Now, knowing the fault and where to fix it, we can use this knowledge to guide the student not only in fixing the issue, but step-by-step in understanding the defect, identifying the potential fix locations, and eventually fixing the problem—similar to how a human tutor would guide the student along the debugging process.
+We designed and developed an AI-assisted debugger that offers students a simulated interactive debugging experience.
 Students receive help in running tests, using the interactive debugger in identifying the underlying fault in the problem, and in fixing the actual problem.
+As Birillo et al.
+[28] have argued, combining programming education with the use of IDEs helps familiarize students with industrial technologies; hence we envision our concept being deployed within IDEs.
 The Role of LLMs
@@
 LLMs can be deployed in various roles in this concept, e.g., to generate hints and explanations in natural language, as well as core technology for an interactive chatbot.
-Based on the generated artifacts from techniques like spectrum-based fault localization (SBFL) (Wong et al., 2016), static code analysis, symbolic execution, code synthesis and LLM prompting, we can set breakpoints automatically, identify relevant variable values, and highlight them for the students.
-Such a process can also be supported by Socratic questioning (Al-Hossami et al., 2024).
+Based on the generated artifacts from techniques like spectrum-based fault localization (SBFL) [29], static code analysis, symbolic execution, code synthesis and LLM prompting, we can set breakpoints automatically, identify relevant variable values, and highlight them for the students.
+Such a process can also be supported by Socratic questioning [23].
 Workflow
@@
 The student provides a buggy solution and (optionally) a set of custom test cases.
-There are three different use cases: (1) the primary use case is the assisted debugging (discussed in Section 3.1), (2) the test assessment for lecturers, and (3) the test assessment for custom tests from the student.
-The test assessment for lecturers checks whether the given test suite is strong enough to catch errors and perform effective fault localization.
-This can be realized via mutation testing, and additional tests could be generated, e.g., via evolutionary test generation (Fraser and Arcuri, 2011).
-By offering additional features for the lecturer, our goal is also to increase confidence in the overall system, which is crucial for its successful deployment in programming courses.
-The test assessment for customs tests from the student works with the idea that when the student submits custom tests, these can be executed against the reference implementation from the lecturer.
-If they fail on the reference implementation, they are flagged and can also be corrected (because the correct output for the inputs of such test cases could be extracted via the reference implementation).
-To help students, we also can propose new test cases to them, e.g., to explore certain boundary cases or to increase overall code coverage.
-Deployment
-Birillo et al.
-(Birillo et al., 2024b) already argued that combining programming education with the usage of IDEs helps to familiarize students with industrial technologies.
-Building on top of this idea, our concept is envisioned to be developed as an IDE plugin.
-3.1. Assisted Debugging
-Figure 1. Envisioned dialogue between student and AI bot.
-Figure 1 showcases the envisioned interaction between the student and the AI assistant.
-First, the AI assistant can (1) clarify/explain failing test cases and the observed failures.
+There are three different use cases: (1) the primary use case is the assisted debugging, (2) the test assessment for lecturers, and (3) the test assessment for custom tests from the student.
+The test assessment for lecturers checks whether the given test suite is strong enough to catch errors and perform effective fault localization, e.g., via mutation testing.
+The test assessment for custom tests from the student works with the idea that when the student submits custom tests, these can be executed against the reference implementation from the lecturer.
+Failing test cases can be flagged and corrected; further, we can propose new tests, e.g., to explore boundary cases or to increase coverage.
+Assisted Debugging
+Figure 1: Envisioned conversation between student and the AI assistant.
+Figure 1 shows the envisioned interaction between student and AI assistant.
+First, the AI assistant can (1) explain failing tests and the observed failures.
 Next, it can (2) automatically set breakpoints at locations of interest.
@@
 Setting such breakpoints enables the unexperienced student to use the interactive debugger to step through the test execution.
-Based on LLMs, we also can (3) generate hints in natural language explaining why these breakpoints are of interest.
-In cases where a student’s understanding of the problem is far from what is expected, the AI assistant can (4) deliver incremental guidance, e.g., via progressive hints, to ensure that the student understands the problem correctly.
-If the student is still stuck, the AI assistant can (5) partially reveal why a certain test case execution fails.
-Furthermore, we can (6) help the student in the interactive debugging by identifying and highlighting interesting program variables that the student can observe during the debugging session.
-This interaction aims to guide the student to repair their own solution and to have a proper ”learning” experience.
-Figure 2.
-Screenshot from our prototypical implementation.
-The left part shows the code and the automatically set breakpoints.
-The right part shows the feature selection and the dialogue with the AI-based chatbot.
-Figure 3. Participants’ programming/debugging experience.
-Figure 4. Participants’ assessment of the provided features.
-4. Pilot User Study
-To provide first insights towards our research goals, we executed a pilot user study. We investigate the practicality of our assisted debugging concept, as well as the usability of our implementation.
-4.1. Implementation
-As a study prototype, we realized the Simulated Interactive Debugging concept as a Visual Studio Code IDE extension.
+Based on LLMs, we also can (3) generate hints in natural language explaining these breakpoints.
+In cases where a student’s understanding of the problem is far from what is expected, the AI assistant can (4) deliver incremental guidance, e.g., via progressive hints.
+If the student is still stuck, the AI assistant can (5) partially reveal why a certain test fails.
+Furthermore, we can (6) help the student by highlighting interesting variables that the student can observe during debugging.
+This interaction aims to guide the student to repair their solution and to have a proper learning experience.
+IV Pilot User Study
+IV-A User Study Setup
+Prototype
+We realize our study prototype as a VS Code IDE extension.
 Our concept is language-agnostic; however, due to the programming courses at our institution, we decided to first focus on supporting Python.
-It provides two core features for the assisted debugger: the automatic setting of breakpoints (based on SBFL with FauxPy (Rezaalipour and Furia, 2024)) and a chat interface providing tailored debugging hints.
-Each feature can be triggered via a VSCode command.
-The chat interface connects the participant with an LLM-powered chatbot.
-The user can choose from two modes: the Generate Hints mode, which only allows the student to get step-by-step hints, and the Interactive Debugging Guidance mode, which allows the student to engage in a dialogue with the LLM in a controlled way.
-The student can discuss code issues and query about how to proceed with debugging or improving the code.
-Figure 2 shows an annotated screenshot of our implementation.
-As underlying LLM, we used OpenAI gpt-3.5-turbo.
-Note that in the implementation of our current prototype, we omitted the support for the lecturer’s and the student’s test assessment because our primary interest was in the feedback for the actual debugging.
-4.2. Study Setup
+It provides two core features for the assisted debugger: the automatic setting of breakpoints (based on SBFL with FauxPy [30]) and a chat interface providing tailored debugging hints.
+The chat interface connects the participant with an LLM-powered chatbot using OpenAI gpt-3.5-turbo.
+Note that in the implementation of our current prototype, we focused only on the debugging use case.
+Figure 2: Participants’ assessment of the provided features.
 Experiment Structure
 The study started with a short briefing, the IRB signing, and the completion of a profiling survey.
-Then, we spent 15 minutes for providing an introduction to the topic of debugging and for demonstrating the features of our tool.
+Then, we provided a 15-minute introduction into debugging and for demonstrating the tool’s features.
 Afterwards, the participants had 40 minutes to solve two coding tasks.
-The tasks originate from LeetCode and represent programming tasks of medium difficulty that should not overwhelm or underwhelm the students.
-Both were provided with pre-existing buggy solutions, along with corresponding passing and failing test cases.
-In the first task, named maxOperations (lee, 2025b), the participants have to fix an algorithm that is supposed to calculate the maximum number of times, to pick two numbers from a given array whose sum equals the parameter $k$.
-In the second task, named longestOnes (lee, 2025a), the participants have to fix the algorithm that is supposed to compute the maximum number of consecutive 1’s in a binary array if one can flip at most $k$ 0’s.
-The participants had first to understand the task’s description, then find the errors in the provided solution, and finally provide a modified implementation that passes all test cases.
-To remember the available assisted debugging features, we also provided each participant with a handout describing a possible workflow to solve the tasks (see replication artifact [^1] Please contact the corresponding author to request the replication artifact.).
-We informed the participants that the chatbot’s hints might not be entirely accurate.
-In the end, the participants were asked to complete a usability questionnaire (Brooke, 1996) and a post-task survey.
+The tasks originate from LeetCode and are of medium difficulty that should not overwhelm or underwhelm the students.
+Both were provided with pre-existing buggy solutions, along with corresponding passing and failing tests.
+In the first task, named maxOperations [31], the participants have to fix an algorithm that is supposed to calculate the maximum number of times to pick two numbers from a given array whose sum equals the parameter $k$.
+In the second task, named longestOnes [32], the participants have to fix the algorithm that is supposed to compute the maximum number of consecutive 1’s in a binary array if one can flip at most $k$ 0’s.
+The participants first had to understand the task’s description, find the errors in the provided solution, and finally modify the program to pass all tests.
+We provided a handout with a possible workflow to solve the tasks with the available assisted debugging features (see artifact).
+We informed the participants that the generated hints might not be entirely accurate.
+In the end, the participants were asked to complete a usability questionnaire [33] and a post-task survey.
 The study was timed for 70 minutes, and the participants were compensated with 30 SGD.
@@
 For our pilot study, we recruited eight first-year undergraduate CS students.
-Based on our profiling survey results, all of them are beginners in the field of programming, and hence, are in our target group.
-Figure 3 illustrates the self-classified experience levels of our participants: Most of them have programming experience of less than two years and can only write basic Python programs.
+Based on our profiling survey results, all of them are beginners in the field of programming, and hence, are in our target group: most of them have programming experience of less than two years and can only write basic Python programs.
 6/8 participants mentioned that they can debug basic errors but struggle with complex ones.
 At the same time, 6/8 have never heard of or never used an interactive debugger.
-Those who understand the concept of an interactive debugger still prefer using print/log statements to debug their code.
-4.3. Results & Discussion
-In the following, we discuss the findings from our user study regarding the feedback on the existing features of the assisted debugger, the overall usability of our prototype, the participant’s confidence of manually setting breakpoints, and additionally desired features.
+Those who understand the concept of interactive debuggers still prefer using print/log statements.
+IV-B Results & Discussion
 Existing Features
@@
 The progressive hints helped the students to identify where their code has a fault but did not reveal the solution.
-Instead, it attempted to guide the students to find the solution on their own to foster a learning effect.
+Instead, it tries to guide the students to find the solution on their own to foster a learning effect.
 However, the participants perceived this feature as often too vague or too general.
 For most participants, the test error explanation helped to analyze the coding issue, while the feature itself could be made more interactively, e.g., allowing for different/alternative explanations.
-The explanation of the automatically set breakpoints helped most students to understand the wrong parts.
+The explanation of the automatically set breakpoints helped most students to understand their faults.
 Therefore, we conclude that it is a good addition to the automatic setting of breakpoints.
@@
 One participant summarized its effectiveness as follows:
-”It did get the breakpoint in the right place, which was a good start in debugging.”
+“It did get the breakpoint in the right place, which was a good start in debugging.”
 One of the first hurdles of conventional interactive debugging is to set breakpoints.
-Our concept of simulated interactive debugging helps them to get started, and therefore, guides them along a deeper understanding of their coding errors and strives to develop debugging skills.
+Our concept of simulated interactive debugging helps the students to get started, and therefore, guides them along a deeper understanding of their coding errors and strives to develop debugging skills.
 The following quote from one of our participants about the interactive debugging and chatting highlights another aspect:
-”It helped me brainstorm while guiding me and
-not letting me fully rely on AI”
+“It helped me brainstorm while guiding me and
+not letting me fully rely on A”
 Our goal is to guide students to learn debugging but also the usage of AI-based tools like ChatGPT.
-By constructing a chatbot that uses LLMs, we can control the prompts and the contents of responses, and hence, still foster a learning experience for the students, e.g., by filtering direct solutions.
+By constructing an LLM-based chatbot, we can control the prompts and the content of responses, and hence, still foster a learning experience for the students, e.g., by filtering direct solutions.
 Usability
-Using the standardized system usability survey by Brooke (Brooke, 1996), we can conclude that overall, the participants are satisfied with the prototype’s usability.
+Using the standardized system usability survey by Brooke [33], we can conclude that overall, the participants are satisfied with the prototype’s usability.
 The average System Usability Scale (SUS) is around 65 (out of 100), providing a good basis for our future work.
 For example, 7/8 participants would like to use the system frequently, and 5/8 agreed that the system is easy to use.
-These findings confirm that our efforts are heading in the right direction but also leave space for further improvement.
 Debugging Process
@@
 The responses indicated mixed confidence with a tendency to lower confidence: 1/8 participants strongly agreed to be confident, 2/8 agreed, 2/8 noted a neutral opinion, and 3/8 disagreed.
-Afterwards, we asked them to set breakpoints for a given code snippet (see Listing 1).
-The interesting lines are 6 to 7 because the handling of None values is missing.
-Listing 1: Code snippet for setting a manual breakpoints.
-1# Calculate the average of a list of
-2# grades, if None, you should continue
-3# to the next grade
-4def calculate_average (grades):
-5 total = 0
-6 for grade in grades:
-7 total += grade
-8 average = total / len (grades)
-9 return average
-10
-11# Test the function
-12grades = [85, 90, 78, None, 92, None]
-13average_grade = calculate_average(grades)
-14print(f”Theaverageis:{average_grade}”)
-6/8 participants would set breakpoints either at lines 6 or 7, or both, which shows that actually most of the participants have a correct intuition about where to set the breakpoint for further inspection.
+Afterwards, we asked them to set breakpoints for a given buggy code snippet that calculates the average for a given list of grades without catering for None values in the list.
+6/8 participants would set breakpoints at the correct lines, which shows that actually most of the participants have a correct intuition about where to set the breakpoint for further inspection.
 While we see good results for this relatively simple example, we conclude that due to overall low confidence, the automated setting of breakpoints would still be an interesting feature.
@@
 The majority (6/8) mentioned that our Simulated Interactive Debugging tool helped them to debug the coding issue and guided them systematically.
-However, one participant mentioned that interactive debugging could be more time-consuming.
+However, one of the participants mentioned that the interactive debugging process could be more time-consuming, though.
 The following two quotes are representative responses:
-”Normally I would just trial and error until I got it right, but this tool allowed me to systematically identify the problems in the code step by step.” ”My usual debugging process includes using print statements and doing everything myself but sometimes I just get tired and ask ChatGPT.
+“Normally I would just trial and error until I got it right, but this tool allowed me to systematically identify the problems in the code step by step.” “My usual debugging process includes using print statements and doing everything myself but sometimes I just get tired and ask ChatGPT.
 Simulated Interactive Debugging helped me think while guiding me through the right process.”
 Additional Features
-At the end of the post-task survey, we asked the participants about any additional features they would like to see in the Simulated Interactive Debugger tool that have not been implemented yet.
+At the end of the post-task survey, we asked the participants about any additional features they would like to see in the Simulated Interactive Debugger.
 6/8 participants indicated that further hints about what and how to change would be helpful.
-For example, when the students really struggle, the chatbot could provide more detailed explanations and propose changes that the student could select.
-Additionally, one participant mentioned that the user interface could be improved to enable easier switching between the debugging information from the assisted debugger and the actual interactive debugging tool in the IDE.
-Lastly, one participant highlighted that integrating some automated test case generation would be helpful.
-For example, such test cases could be provided if the students cannot find an appropriate input for debugging.
-4.4. Threats to Validity
-Since we only performed a first pilot study with eight CS students, our results may not generalize.
-We focused our user study on a small number of students to receive focused feedback for the next development phase.
-Nevertheless, the results of our profiling survey showed that the participants are in our target, i.e., beginners in programming with no or only very limited experience with (interactive) debugging.
-For the next cycle of user studies, we will attempt to reach a larger set of participants.
-In order to mitigate the threat of manual errors in our analysis, we ensured that all results were analyzed and agreed upon by two authors independently.
-5. Conclusion and Future Work
+For example, when a student really struggles, the chatbot could provide more detailed explanations and propose changes that the student could select.
+One participant mentioned that the UI could be improved to allow easier switching between the assisted debugger and the interactive debugging tool in the IDE.
+Lastly, one participant highlighted that automated test generation would be helpful, e.g., if students cannot find an appropriate input for debugging.
+IV-C Threats to Validity
+We performed a first, small pilot study with eight CS students, i.e., our results may not generalize.
+We focused on a small number of students to receive focused feedback for the next development phase.
+The results of our profiling survey showed that the participants are in our target, i.e., beginners in programming with no or very limited (interactive) debugging experience.
+For the next cycle of user studies, we will try to reach a larger set of participants.
+To mitigate the threat of manual errors in our analysis, we ensured that all results were analyzed and agreed upon by two authors independently.
+V Conclusion and Future Work
 We proposed the concept of Simulated Interactive Debugging as a first step to automate the teaching of debugging skills.
-Incorporating such techniques in the CS curriculum will be essential to teaching debugging beyond ad-hoc approaches and guiding students in using AI-based programming tools.
-In future, we plan to improve our interactive chatbot with more guidance, e.g., by using a state machine-driven approach, similar to Bouzenia et al.
-(Bouzenia et al., 2024).
-We also will explore the integration of several state-of-the-art APR techniques, implement the test assessment features for lecturers and students, and investigate how we can automatically help students to generate and improve their test suites.
+Incorporating such techniques in the CS curriculum will be essential to teach debugging beyond ad-hoc approaches and guide students in using AI-based programming tools.
+In future, we plan to incorporate more active guidance, e.g., by using a state machine-driven approach, similar to Bouzenia et al.
+[34].
+We will also integrate state-of-the-art APR techniques and implement the test assessment and generation features.
+Data Availability
+Our artifact with the prototype and the study artifacts is available: https://doi.org/10.6084/m9.figshare.28202336
+Acknowledgment
+Part of this work was supported by the Ministry of Education, Singapore, under the Tertiary Research Fund (MOE-TRF) Grant No.
+MOE2023-TRF-034.
+We gratefully acknowledge this support, which made this research possible.